OpenStudy (anonymous):

Use substitution to find the value of y for this set of equations. 12x - 5y = 30 y = 2x - 6

OpenStudy (*insert name here*):

Sub in y: \[\Large 12x-5(2x-6)=30\] Now solve for x.

OpenStudy (anonymous):

To use substitution, we must first isolate x or y in either of the equations. What is awesome here, is y is isolated already in the 2nd equation!!! y = 2x - 6 So! Let's just plug in the 2x - 6 into the y of the first equation. 12x - 5y = 30; And we know y = 2x - 6 So,\[12x - 5(2x-6) = 30\] Now solve for x!

OpenStudy (anonymous):

Do you know what to do now? @MyShoeGameIsMagic

OpenStudy (anonymous):

yes thanks!

OpenStudy (anonymous):

Of course bud!